OpenStudy (anonymous):

I GIVE MEDALS!!!!!! Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6). A.) y = 1/3x + 7 B.) y = −3x + 3 C.) y = 1/3x + 3 D.) y = −3x + 7

OpenStudy (anonymous):

look at da problem close ly

OpenStudy (anonymous):

what is the slope for the given line ? Can you find that ?

OpenStudy (anonymous):

In y = mx + b form, the slope is in the m position. Now look at your line, y = -3x + 12, can you find the slope ?

OpenStudy (anonymous):

good job @kelliegirl33

OpenStudy (anonymous):

I am trying to explain, but is he even paying attention ??

OpenStudy (anonymous):

I will not explain anymore unless turtleluver says something

OpenStudy (anonymous):

it is a girl and probley not lol

OpenStudy (anonymous):

oh...sorry....she

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

:) lol

OpenStudy (anonymous):

wut u doing

OpenStudy (anonymous):

well...when she wants to participate, she can tag me....I can see my notifications now...HOORAY

OpenStudy (anonymous):

hey can u explain it 2 me plz?

OpenStudy (anonymous):

why yes I can :) do you know the slope of the line given ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

In y = mx + b form, the slope is in the m position. So in y = -3x + 12, what is the slope ? What number is in the m position ?

OpenStudy (anonymous):

do you know ?

OpenStudy (anonymous):

ok.....I am just gonna do this one... y = -3x + 12....the slope is -3. A parallel line will have the same slope y = mx + b slope(m) = -3 (-1,6) x = -1 and y = 6 (we already know the slope, so we just need the y intercept, which is in the b position, so we have to find b. now we sub y = mx + b 6 = -3(-1) + b 6 = 3 + b 6 - 3 = b 3 = b our parallel line is : y = -3x + 3

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

good to hear :)

OpenStudy (anonymous):

lol sorry no response i had to take care of meah dog

OpenStudy (anonymous):

there is also another way to do this....do you want to know how ?

OpenStudy (anonymous):

if you were asked to put this in point slope form, rather then slope intercept form, you would have to do it another way

OpenStudy (anonymous):

um.... sure

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

point slope form is : y - y1 = m(x - x1) slope(m) = -3 (-1,6) x1 = -1 and y1 = 6 now we sub y - 6 = -3(x - (-1) y - 6 = -3(x + 1) <== this would be your point slope form answer we can continue to slope intercept form from here y - 6 = -3x - 3 y = -3x - 3 + 6 y = -3x + 3 <== this is your slope intercept form

OpenStudy (anonymous):

ohhh ok thx CX

OpenStudy (anonymous):

so if there asking for slope intercept, you can do y = mx + b, but if their asking for point slope form, you have to do y - y1 = m(x - x1)

OpenStudy (anonymous):

kk thx

OpenStudy (anonymous):

no problem :)